# Estimating Machine Translation Difficulty

### v1

@@ Abstract
Machine translation quality has began achieving near-perfect translations in some setups. These high-quality outputs make it difficult to distinguish between state-of-the-art models and to identify areas for future improvement. Automatically identifying texts where machine translation systems struggle holds promise for developing more discriminative evaluations and guiding future research.
We formalize the task of translation difficulty estimation, defining a text’s difficulty based on the expected quality of its translations. We introduce a new metric to evaluate difficulty estimators and use it to assess both baselines and novel approaches. Finally, we demonstrate the practical utility of difficulty estimators by using them to construct more challenging machine translation benchmarks. Our results show that dedicated models (dubbed Sentinel-src) outperform both heuristic-based methods (e.g. word rarity or syntactic complexity) and LLM-as-a-judge approaches. We release two improved models for difficulty estimation, Sentinel-src-24 and Sentinel-src-25, which can be used to scan large collections of texts and select those most likely to challenge contemporary machine translation systems.
@@ 1 Introduction
Evaluating models also benefits from estimates of sample difficulty, as too easy or too difficult benchmarks fail to effectively differentiate between models Lalor et al. (2018); Rodriguez et al. (2021). This issue is particularly relevant in Machine Translation (MT), with recent state-of-the-art models obtaining near-perfect scores and performing close to the human level Kocmi et al. (2024a). With easy test sets, practitioners might struggle to differentiate between top-performing models and assess whether there is headroom for further model improvement. Additionally, while the MT Test Suites subtask of WMT (Kocmi et al., 2024a) targets specific complex translation phenomena, no systematic investigation of the broader concept of general translation difficulty has been carried out.
@@
We find that approaches such as word rarity, syntactic complexity, or even LLM-as-a-Judge underperform dedicated solutions in capturing translation difficulty. Specifically, we find Sentinel-src– a model trained to predict the expected translation quality of a given text based solely on the source text itself Perrella et al. (2024) – to outperform other methods at estimating translation difficulty. Therefore, we train two improved versions, called Sentinel-src-24 and Sentinel-src-25, and release them publicly. [^1] Models: hf.co/collections/Prosho/translation-difficulty-estimators-6816665c008e1d22426eb6c4
Code: github.com/zouharvi/translation-difficulty-estimation
@@ 2 Related Work
To implement a Curriculum Learning training schedule, Kocmi and Bojar (2017) estimated sample difficulty based on sentence length, word rarity, and the number of coordinating conjunctions in the text. Similarly, Platanios et al. (2019) used sentence length and rarity as proxies for difficulty. Beyond these linguistically-motivated criteria, Zhang et al. (2018) and Liu et al. (2020) predicted translation difficulty using the confidence and other intrinsics of the translation model in generating the text. Almeida (2017) treated difficulty estimation as a binary classification task, though uses features also from the target text, making it closer to quality estimation. Zhan et al. (2021b) used artificial crowd with automated metrics and discovered that long segments, and low-frequency and proper nouns are the most challenging to machine translate. Finally, Zhan et al. (2021a) estimated text’s difficulty using the embedding similarity between its tokens and those of its translations.
Close to our work, Don-Yehiya et al. (2022) defined the PreQuEL task as predicting the quality of a given text’s translation before the translation is generated. However, they adopted the evaluation of the WMT 2020 Quality Estimation Shared Task Specia et al. (2020), which was designed for quality estimation rather than for assessing difficulty estimators. Furthermore, their test set included only two language directions, with all translations produced by the same MT model. Additionally, they did not explore the broader space of difficulty estimators or investigate their use in constructing challenging benchmarks.
@@ 3 The Difficulty Estimation Task
The difficulty of translating a given text can depend on multiple factors. A text may be challenging, for example, due to its length, syntactic complexity, idiomatic language, or the presence of rare or specialized vocabulary. Some of the aspects that affect translation difficulty may even depend on the translation direction, meaning that the same source text might be more difficult to translate into one language than into another. Moreover, translation difficulty might not be uniform across translators. It can vary with the translator’s cultural background and linguistic familiarity, in the case of human translators, or based on factors such as the number of parameters, training data, and model architecture, in the case of machine translation models.
Given these considerations, we avoid defining translation difficulty in absolute terms, as such a definition may not generalize well. Instead, we define difficulty relative to a given target language and to the accuracy of a particular translator, whether human or automatic. More specifically, given a text $x$, a translator $m$, and a target language $l$, we assign a difficulty score $d_{m,l}(x)$ equal to the quality score assigned to a translation of $x$ into language $l$ produced by $m$. Lower scores indicate a lower translation quality and, therefore, greater difficulty associated with the source text.
As an example, suppose we have two texts, $x_{1}$ and $x_{2}$, and their respective translations $t_{1}$ and $t_{2}$ into language $l$, both produced by model $m$. A human rater evaluates these translations on a scale from $1$ to $100$, assigning a score of $60$ to the first and $90$ to the second. Then, $d_{m,l}(x_{1})=60$ and $d_{m,l}(x_{2})=90$. Since $d_{m,l}(x_{1})<d_{m,l}(x_{2})$, then $x_{1}$ is more difficult to translate into $l$ than $x_{2}$, for translator $m$. Importantly, the lower the score $d$, the higher the difficulty and vice versa.
@@
Given a source text $x$, a translator $m$, and a target language $l$, Difficulty Estimation is the task of automatically predicting $d_{m,l}(x)$. Different from Quality Estimation, difficulty estimation models do not have access to the translations whose quality is being estimated. Indeed, difficulty estimation can be seen as the task of estimating the expected quality of a given text’s translation.
@@
We evaluate difficulty estimation methods according to their ability to rank texts based on difficulty. Consider a collection of texts $\mathcal{X}=x_{1},x_{2},\dots,$ $x_{N}$, a collection of target languages $\mathcal{L}=l_{1},$ $l_{2},\dots,l_{L}$, and a collection of models [^2] For brevity, we use “model” to refer both to to human translators and automatic models. translating into language $l$: $\mathcal{M}_{l}=m_{1},m_{2},\dots,m_{M_{l}}$. Let us also define the vector of ground-truth difficulty scores for model $m$ and language $l$ as $D_{m,l}=d_{m,l}(x_{1}),$ $d_{m,l}(x_{2}),\dots,d_{m,l}(x_{N})$ and the corresponding predictions of a difficulty estimation method as $\hat{D}_{m,l}=\hat{d}_{m,l}(x_{1}),$ $\hat{d}_{m,l}(x_{2}),$ $\dots,$ $\hat{d}_{m,l}(x_{n})$. We measure the translation Difficulty Estimation Correlation (DEC) by averaging the Kendall’s rank correlation coefficients $\tau_{b}$ across translators and languages:
@@
However, we argue that these features are precisely what defines a text’s translation difficulty. Since our goal is to measure translation difficulty, we define DEC using Group-by-System. This method correlates human and metric assessments for translations of different source texts that were produced by the same MT system, and then averages the correlations across MT systems. By holding the MT system constant, this evaluation directly measures a metric’s ability to identify which source texts were more challenging for that system to translate.
@@ 4.1 Heuristic-based estimators
  • Text length as the number of words in a text.
@@
  • Syntactic complexity is approximated as the height of the dependency tree of a text.
@@ 4.2 Learned estimators
  • PreCOMET is a suite of source-based regressors based on XLM-RoBERTa Conneau et al. (2020) that predict the usefulness of a sample for evaluation (Zouhar et al., 2025b). PreCOMETdiversity prioritizes samples likely to elicit diverse machine translation outputs, while PreCOMETdifficulty estimates difficulty as defined by item response theory (Santor and Ramsay, 1998).
  • Sentinel-src metrics are XLM-RoBERTa-based models trained to predict the quality of a text’s translations without taking the translation as input. The objective is to learn spurious correlations between the source text and the translation quality scores Perrella et al. (2024).
@@ 4.3 LLM-as-a-Judge
LLM-as-a-Judge approaches have seen wide adoption across a range of applications Zheng et al. (2023); Bavaresco et al. (2024). In this work, we investigate the effectiveness of the LLM-as-a-Judge paradigm for the task of difficulty estimation, using GPT-4o OpenAI et al. (2024) and CommandA Cohere et al. (2025). We prompt these models to determine the proficiency level required to translate a given text, optionally providing information about the target language, and return a scalar score between 0 and 120 indicating the difficulty level of the given text. See the prompts in Figure˜7.
@@ 4.4 Crowd-based Estimators
Artificial crowd-based methods first translate a source text and then use reference-less MT metrics to estimate the quality of the resulting translations. [^3] Reference-less MT metrics estimate the quality of a translation by comparing it only to its source text, without requiring reference translations. Specifically, we translate the source texts from the test set using a diverse set of models to ensure variety in architecture and size: three LLM-based machine translation models (Gemma-3-27B-IT, Qwen2.5-72B-IT, CommandA) and one standard encoder-decoder machine translation model (NLLB-moe-54B). For the evaluation step, we employ two two state-of-the-art, reference-less MT metrics: XCOMET-QE-XXL Guerreiro et al. (2024a) and MetricX-24-Hybrid-QE-XXL Juraska et al. (2024), hereafter referred to as XCOMET and MetricX, respectively. The final difficulty score for each source text is the average quality score assigned to its translations by one of these metrics. This approach is inspired by the artificial crowd methods for efficient subset selection proposed by Zouhar et al. (2025b).
@@
To establish a performance upper bound for our artificial crowd-based estimators, we also define True Crowd estimators. They use XCOMET or MetricX to score the translations from the actual systems for which we are measuring difficulty, that is, the systems whose translations were annotated by humans for quality. The True Crowd approach is therefore equivalent to standard quality estimation, and its performance depends entirely on the accuracy of the underlying reference-less metrics.
@@ 5.1 Experimental Setup
Table 1: Difficulty Estimation Correlation (DEC) achieved by each method. We categorize the methods based on the type of information they have access to. Text-only estimators, such as the heuristic and learned ones, rely solely on the source text whose difficulty is being estimated. Instead, some methods also incorporate information of the target language (Lang.) or of the specific translator who produced the translation (Trans.).
@@ 5.2 Results
We show the results in Table˜1, with methods grouped according to their categories from Section˜4. We also mark each method with the information it uses (i.e., model or language), as detailed in the caption of Table˜1, and include three distinct oracles to provide the reader with upper-bound performance values. The definition of Oracles can be found in Appendix˜E.
@@
These estimators base their predictions only on the input text. Consequently, the difficulty scores they assign to each text are the same across all target languages and models. [^5] I.e., $\forall m_{1},m_{2}{\in}\mathcal{M},l_{1},l_{2}{\in}\mathcal{L}:\,\hat{d}_{m_{1},l_{1}}(x)=\hat{d}_{m_{2},l_{2}}(x)$. Within this group, all learned estimators outperform the heuristic-based ones. Furthermore, Sentinel-src-24 achieves the highest difficulty estimation correlation, also higher than Sentinel-src from Perrella et al. (2024), highlighting the effectiveness of our re-training.
@@
As expected, the True Crowd methods, which utilize ground truth translations and thus serve as an upper bound, yield the highest correlation. Since their performance depends entirely on the employed reference-less metrics, this result also demonstrates that XCOMET outperforms MetricX on this task by a noticeable margin.
@@ Discussion.
  • Heuristic-based estimators, which were commonly used in previous works, are outperformed by most other methods.
@@
  • Learned methods, i.e., models explicitly trained to predict the difficulty of a given text, outperform most other approaches.
  • Sentinel-src-24 is a state-of-the-art difficulty estimator, matched only by Artificial Crowd estimators that require much larger computational efforts.
Based on these results, the next section examines the ability of the top-performing estimators from each category in the downstream task of creating difficult benchmarks. To ensure fairness, we exclude certain artificial crowd-based estimators. These methods use translations generated by multiple MT models to estimate text difficulty, which would unfairly bias the resulting test set against the MT models used in this process. Specifically, we only include the top-performing external artificial crowd-based estimator, assuming the generated test data will be used to evaluate models other than those involved in its construction.
@@ 5.3 Comparing Human and Machine Translation Difficulty
In this section, we examine whether the texts that models find difficult to translate are also challenging for humans. To do this, we use the difficulty scores $d_{m,l}$ assigned to each source text in the WMT 24 test sets, varying $m$ across human translators and MT models. We measure the Kendall’s $\tau_{b}$ between the scores of all pairs of translators, averaging them across all language directions. For consistency, we restrict the analysis to the models and human translators for which we have annotated translations for all language directions, namely, one human translator and the following four models: Unbabel-Tower70B Alves et al. (2024), IOL-Research Zhang (2024), Claude-3.5, and GPT-4 OpenAI et al. (2024).
@@
Figure 1: Distribution of human scores assigned to the translations of the texts included in the WMT 2024 test set (Kocmi et al., 2024a). We report the scores of all models (ALL), the scores of the top-performing model for each language (Top system), and the scores of the best translation for each input text (Top translation). The chosen bin width is 15 ESA points.

### v2

@@ Abstract
Machine translation quality has steadily improved over the years, achieving near-perfect translations in recent benchmarks. These high-quality outputs make it difficult to distinguish between state-of-the-art models and to identify areas for future improvement. In this context, automatically identifying texts where machine translation systems struggle holds promise for developing more discriminative evaluations and guiding future research.
In this work, we address this gap by formalizing the task of translation difficulty estimation, defining a text’s difficulty based on the expected quality of its translations. We introduce a new metric to evaluate difficulty estimators and use it to assess both baselines and novel approaches. Finally, we demonstrate the practical utility of difficulty estimators by using them to construct more challenging benchmarks for machine translation. Our results show that dedicated models outperform both heuristic-based methods and LLM-as-a-judge approaches, with Sentinel-src achieving the best performance. Thus, we release two improved models for difficulty estimation, Sentinel-src-24 and Sentinel-src-25, which can be used to scan large collections of texts and select those most likely to challenge contemporary machine translation systems.
@@ 1 Introduction
Evaluating models also benefits from estimates of sample difficulty, as too easy or too difficult benchmarks fail to effectively differentiate between models Lalor et al. (2018); Rodriguez et al. (2021). This issue is particularly relevant in Machine Translation (MT), with recent state-of-the-art models obtaining near-perfect scores and performing close to the human level Kocmi et al. (2024a); Proietti et al. (2025). With easy test sets, practitioners might struggle to differentiate between top-performing models and assess whether there is headroom for further model improvement. Additionally, while the MT Test Suites subtask of WMT (Kocmi et al., 2024a) targets specific complex translation phenomena, no systematic investigation of the broader concept of general translation difficulty has been carried out.
@@
We find that approaches such as word rarity, syntactic complexity, or even LLM-as-a-Judge underperform dedicated solutions in capturing translation difficulty. Specifically, we find Sentinel-src– a model trained to predict the expected translation quality of a given text based solely on the source text itself Perrella et al. (2024) – to outperform other methods at estimating translation difficulty. Therefore, we train two improved versions, called Sentinel-src-24 and Sentinel-src-25, and release them publicly. [^1] Models: hf.co/collections/Prosho/translation-difficulty-estimators-6816665c008e1d22426eb6c4.
Code: github.com/zouharvi/translation-difficulty-estimation.
@@ 2 Related Work
To implement a Curriculum Learning training schedule, Kocmi and Bojar (2017) estimated sample difficulty based on sentence length, word rarity, and the number of coordinating conjunctions in the text. Similarly, Platanios et al. (2019) used sentence length and rarity as proxies for difficulty. Beyond these linguistically-motivated criteria, Zhang et al. (2018) and Liu et al. (2020) predicted translation difficulty using the confidence and other intrinsics of the translation model in generating the text. Almeida (2017) treated difficulty estimation as a binary classification task, although they also used features from the target text, making it closer to quality estimation. Zhan et al. (2021b) used an artificial crowd-based approach that leverages automatic metrics and discovered that long segments, low-frequency words, and proper nouns are the most challenging to machine translate. Finally, Zhan et al. (2021a) estimated text’s difficulty using the embedding similarity between its tokens and those of its translations.
Closer to our work, Don-Yehiya et al. (2022) defined the PreQuEL task as predicting the quality of a given text’s translation before the translation is generated. However, they adopted the evaluation of the WMT 2020 Quality Estimation Shared Task Specia et al. (2020), which was designed for quality estimation rather than for assessing difficulty estimators. Furthermore, their test set included only two language directions, with all translations produced by the same MT model. Additionally, they did not explore the broader space of difficulty estimators or investigate their use in constructing challenging benchmarks.
@@ 3 The Difficulty Estimation Task
The difficulty of translating a given text can depend on multiple factors. A text may be challenging, for example, due to its length, syntactic complexity, idiomatic language, or the presence of rare or specialized vocabulary. Some aspects that affect translation difficulty may even depend on the translation direction, meaning that the same source text might be more difficult to translate into one language than into another. Moreover, translation difficulty might not be uniform across translators, as it can vary with the translator’s cultural background and linguistic familiarity – in the case of human translators – or based on factors such as the number of parameters, training data, and model architecture – in the case of machine translation models.
Given these considerations, we avoid defining translation difficulty in absolute terms, as such a definition may not generalize well. Instead, we define difficulty relative to a given target language and to the accuracy of a particular translator, whether human or automatic. More specifically, given a text $x$, a model [^2] For brevity, we use “model” to refer both to human translators and automatic models. $m$, and a target language $l$, we assign to $x$ a difficulty score $d_{m,l}(x)$ equal to the quality score assigned to a translation of $x$ into language $l$ produced by $m$. Lower scores indicate a lower translation quality and, therefore, greater difficulty associated with the source text.
As an example, suppose we have two texts, $x_{1}$ and $x_{2}$, and their respective translations $t_{1}$ and $t_{2}$ into language $l$, both produced by model $m$. A human rater evaluates these translations on a scale from $1$ to $100$, assigning a score of $60$ to the first and $90$ to the second. Then, $d_{m,l}(x_{1})=60$ and $d_{m,l}(x_{2})=90$. Since $d_{m,l}(x_{1})<d_{m,l}(x_{2})$, then $x_{1}$ is more difficult to translate into $l$ than $x_{2}$, for model $m$. Importantly, the lower the score $d$, the higher the difficulty and vice versa.
@@
Given a source text $x$, a model $m$, and a target language $l$, Difficulty Estimation is the task of automatically predicting $d_{m,l}(x)$. Different from Quality Estimation, difficulty estimation models do not have access to the translations whose quality is being estimated. Indeed, difficulty estimation can be seen as the task of estimating the expected quality of a given text’s translation.
@@
We evaluate difficulty estimation methods according to their ability to rank texts based on difficulty. Consider a collection of texts $\mathcal{X}=x_{1},x_{2},\dots,$ $x_{N}$, a collection of target languages $\mathcal{L}=l_{1},$ $l_{2},\dots,l_{L}$, and a collection of models translating into language $l$: $\mathcal{M}_{l}=m_{1},m_{2},\dots,m_{M_{l}}$. Let us also define the vector of ground-truth difficulty scores for model $m$ and language $l$ as $D_{m,l}=d_{m,l}(x_{1}),$ $d_{m,l}(x_{2}),\dots,d_{m,l}(x_{N})$ and the corresponding predictions of a difficulty estimation method as $\hat{D}_{m,l}=\hat{d}_{m,l}(x_{1}),$ $\hat{d}_{m,l}(x_{2}),$ $\dots,$ $\hat{d}_{m,l}(x_{n})$. We measure the translation Difficulty Estimation Correlation (DEC) by averaging the Kendall’s rank correlation coefficients $\tau_{b}$ across models and languages:
@@
However, we argue that these features are precisely what define a text’s translation difficulty. Since our goal is to measure translation difficulty, we define DEC using Group-by-System. This method computes the correlation between human and metric assessments for translations of different source texts that were produced by the same MT system, and then averages the correlations across MT systems. By holding the MT system constant, this evaluation directly measures a metric’s ability to identify which source texts were more challenging for that system to translate.
@@ 4.1 Heuristic-based estimators
  • Text length is the number of tokens in a text.
@@
  • Syntactic complexity is approximated as the height of the dependency tree associated with a text.
@@ 4.2 Learned estimators
  • PreCOMET is a suite of source-based regressors based on XLM-RoBERTa Conneau et al. (2020) that predict the usefulness of a sample for evaluation (Zouhar et al., 2025b). Specifically, PreCOMETdiversity prioritizes samples likely to elicit diverse machine translation outputs, while PreCOMETdifficulty estimates difficulty as defined by item response theory (Santor and Ramsay, 1998).
  • Sentinel-src metric is a regression model based on XLM-RoBERTa. Perrella et al. (2024) trained Sentinel-src to estimate translation quality from the source alone – i.e., without accessing the candidate translation – with the goal of learning spurious correlations between features of the source texts and translation quality scores.
@@ 4.3 LLM-as-a-Judge
LLM-as-a-Judge approaches have seen wide adoption across a range of applications Zheng et al. (2023); Bavaresco et al. (2024). In this work, we investigate the effectiveness of the LLM-as-a-Judge paradigm for the task of difficulty estimation, using GPT-4o OpenAI (2024) and CommandA Cohere Team (2025). We prompt these models to determine the proficiency level required to translate a given text, optionally providing information about the target language, and return a scalar score between 0 and 120 indicating the difficulty level of the given text. See the prompts in Figure˜8.
@@ 4.4 Crowd-based Estimators
Artificial crowd-based methods first translate a source text and then use reference-less MT metrics to estimate the quality of the resulting translations. [^3] Reference-less MT metrics estimate the quality of a translation by comparing it only to its source text, without requiring reference translations. Specifically, we translate the source texts from the test set using a diverse set of models to ensure variety in architecture and size: three instruction-tuned LLMs (Gemma-3-27B-IT, Qwen2.5-72B-IT, CommandA) and one standard encoder-decoder machine translation model (NLLB-moe-54B). For the evaluation step, we employ two state-of-the-art, reference-less MT metrics: XCOMET-QE-XXL Guerreiro et al. (2024a) and MetricX-24-Hybrid-QE-XXL Juraska et al. (2024), hereafter referred to as XCOMET and MetricX, respectively. The final difficulty score for each source text is the average quality score assigned to its translations by one of these metrics. This approach is inspired by the artificial crowd methods for efficient subset selection proposed by Zouhar et al. (2025b).
@@
To establish a performance upper bound for Artificial Crowd estimators, we also define True Crowd estimators. Unlike Artificial Crowd, True Crowd estimators use XCOMET and Metric to score the translations produced by the actual systems whose difficulty we aim to measure – namely, the translations that constitute the WMT24 test set used in our experiments.
Since they rely on the “ground-truth” translations, True Crowd estimators are effectively equivalent to quality estimators. Thus, they are not proper difficulty estimators; we employ them solely to report an upper bound on the performance of Artificial Crowd estimators.
@@ 5.1 Experimental Setup
Table 1: Difficulty Estimation Correlation (DEC) achieved by each method. We categorize the methods based on the type of information they have access to. Text-only estimators, such as the heuristic and learned ones, rely solely on the source text whose difficulty is being estimated. Instead, some methods also incorporate information of the target language (Lang.) or of the true translation included in the test set (Trans.).
@@ 5.2 Results
We present the results in Table˜1, with methods organized by category as described in Section˜4. We also mark each method with the information it uses (i.e., true translations or target language), as detailed in the caption of Table˜1, and include three distinct oracles to provide the reader with upper-bound performance values. The definition of Oracles can be found in Appendix˜E.
@@
These estimators base their predictions only on the input text. Consequently, the difficulty scores they assign to each text are the same across all target languages and models. [^5] I.e., $\forall m_{1},m_{2}{\in}\mathcal{M},l_{1},l_{2}{\in}\mathcal{L}:\,\hat{d}_{m_{1},l_{1}}(x)=\hat{d}_{m_{2},l_{2}}(x)$. Within this group, all learned estimators outperform the heuristic-based ones. Furthermore, Sentinel-src-24 achieves the highest difficulty estimation correlation overall, also higher than Sentinel-src from Perrella et al. (2024), highlighting the effectiveness of our re-training.
@@
As expected, the True Crowd methods, which utilize ground-truth translations and thus serve as an upper bound, yield the highest correlation. Since their performance depends solely on the employed reference-less metrics, this result also demonstrates that XCOMET outperforms MetricX on this task by a noticeable margin.
@@ 5.3 Discussion
  • Heuristic-based estimators, commonly used in previous works, are outperformed by most other methods.
@@
  • The performance of learned methods – i.e., models explicitly trained to predict text difficulty – is matched only by Artificial Crowd estimators, which are, however, considerably more expensive to operate.
  • Sentinel-src-24 sets a state-of-the-art in difficulty estimation, outperforming all other evaluated estimators. [^6] We intentionally exclude True Crowd estimators, which, as discussed in Section 4.4, are not genuine difficulty estimators and instead serve only as upper bounds for Artificial Crowd.
Based on these results, Section˜6 examines the ability of the top estimators from each category – excluding True Crowd estimators, for the reasons discussed in Section˜4.4 – on the downstream task of constructing difficult benchmarks.
Furthermore, even if Artificial Crowd estimators are included, using them for benchmark creation implicitly assumes that the generated test data will be used to evaluate models other than those involved in the difficulty estimation. In fact, because Artificial Crowd relies on an intermediate translation step, it would bias the resulting test set against the models used in its construction.
Full results with significance testing – including those restricted to the MQM-annotated portion of WMT24 – are provided in Appendix˜D.
@@ 5.4 Comparing Human and Machine Translation Difficulty
In this section, we examine whether the texts that models find difficult to translate are also challenging for humans. To do this, we use the difficulty scores $d_{m,l}$ assigned to each source text in the WMT 24 test sets, varying $m$ across human translators and MT models. We measure the Kendall’s $\tau_{b}$ between the scores of all pairs of translators, averaging them across all language directions. For consistency, we restrict the analysis to the models and human translators for which we have annotated translations for all language directions, namely, one human translator and the following four models: Unbabel-Tower70B Alves et al. (2024), IOL-Research Zhang (2024), Claude-3.5, and GPT-4 OpenAI (2024).
@@
Figure 1: Distribution of human scores assigned to the translations of the texts included in the WMT 2024 test set (Kocmi et al., 2024a). We report the scores of all models (ALL), the scores of the top-performing model for each language (Top model), and the scores of the best translation for each input text (Top translation). The chosen bin width is 15 ESA points.